OpenStudy (anonymous):

Abe's car can run 32 for miles with one gallon of gasoline. He just filled up his 14-gallon tank with gasoline. If he drives his car 40 miles a day, how many full days can be go before needing to get gas again?

OpenStudy (shane_b):

Start with this: If he gets 32 mpg and has 14 gallons of gas, how many total miles can he go?

OpenStudy (anonymous):

Total miles that Abe can run using 14-gallons of gasoline=32*14 Number of miles run per day=40 Number of days =Total miles that Abe can run using 14-gallons of gasoline/Number of miles run per day =32*14/40 Can u do now?